# Evolving Agents: Interactive Simulation of Dynamic and Diverse Human Personalities

### v1

@@ 6.1. Evaluation Procedure
We recruited a total of 30 human evaluators (14 male, 16 female) who signed informed consent forms to participate in our study. Their ages ranged from 18 to 34 years, encompassing a variety of disciplines and professions, with good reading and analytical skills, and all having a bachelor’s degree or higher education.
Before the formal experiment, we conducted a preliminary study. In this pre-experiment, we first selected a 10-day log of one of the agents and presented it through front-end animations on the simulation platform for two human evaluators to assess. During the pre-experiment, we identified the following issues: 1) In practice, we found that human evaluators were overly distracted by images and other irrelevant information, making it difficult for them to focus on the agent’s personality growth and behavioral changes themselves. 2) The amount of textual information from 10 days of evolution was too overwhelming for general evaluators, making it difficult for them to fully understand the complete process of the agent’s changes over the 10 days; 3) The evolution of the agent’s personality and behavior is a highly abstract and generalized simulation of human personality evolution. Because we cannot observe an agent for over a decade or more like we do with humans. Human evaluators need to detach from the normal human perception of time and focus on evaluating the evolution process itself.
Therefore, in the formal experiment, we used text records of an agent’s logs over three consecutive days for user assessment. Additionally, before the experiment, we emphasized to the human evaluators the importance of reducing consideration of the time factor and focusing on the believability of the evolutionary process itself.
In the formal experiment, we first asked human evaluators to spend about an hour reading and taking notes on the logs of three agents and the logs of four ablated versions of one agent. We also asked them to use the think aloud method to communicate their thoughts to the experimenters during the reading process. After completing the reading, we asked them to fill out questionnaires for both the control and ablation experiments based on their notes. During this process, they could continually refer back to the reading materials to ensure the accuracy of their choices. The questionnaires we set up focused on two main aspects: user believability rating, and ablation control experiments.
@@ 6.3. Ablation Experiment
To validate the positive impact of various modules within the Evolving Agents on both personality differentiation and personality evolution in terms of perceptibility and believability, we conducted two degrees of ablation studies on the architecture of the Evolving Agent. Compared to the full Evolving Agents architecture, Group 2 had the Cognitive Feelings, Insight, and Character Growth parts of the Personality system removed, consistent with section 5.1, while Group 3 was further simplified from Group 2 by simplifying the Character Structure, adopting a character description method similar to those used in other existing embodied agent sandbox simulation systems. We asked human evaluators to rank the human-likeness of the growth trajectories of personality and behavior for four experimental groups and to describe in words the reasons for their rankings.
@@
In the ablation experiment, we obtained 30 sets of ranking data from user studies. Based on this data, we calculated the TrueSkill ranking scores for the four experimental groups and computed Cohen’s d values to quantify the strength of differences between groups. In the TrueSkill scoring system, the $\mu$ value represents the average performance level of an experimental group, while the $\sigma$ value measures the uncertainty of the scores. Additionally, to explore the statistical significance of these results, we conducted a Kruskal-Wallis test, a non-parametric alternative to one-way ANOVA, to confirm the overall statistical significance of differences under different experimental groups. To pinpoint the exact sources of differences, namely which specific group comparisons showed significant differences, we used the Dunn test for pairwise comparisons between experimental groups, and p-values were adjusted using the Holm-Bonferroni method. The experimental results are illustrated in a bar chart, with the height of the bars corresponding to the $\mu$  value and the top of the bars marked with error lines, indicating the $\sigma$  value.
@@
The final result shows, Group 1 achieved the best scoring results ($\mu$=29.89; $\sigma$=1.61), followed by Group 2 ($\mu$=23.18; $\sigma$=1.13), and then Group 3 ($\mu$=22.95; $\sigma$=1.15). Comparing the highest-scoring Group 1 with the composite scores of the other three groups, the Cohen’s d value of 5.33 indicates that Group 1 performed significantly better than Groups 2, 3, and 4 combined. The Dunn post-hoc test, both uncorrected and corrected using the Holm-Bonferroni method, showed that the difference between Group 1 and all other groups (Groups 2, 3, and 4) remained significant, while there was no significant difference among Groups 2, 3, and 4.
@@
Among all experimental control groups, Group 4 performed the worst ($\mu$=21.16; $\sigma$=1.18 ). Surprisingly, the human group was perceived by users to have the lowest level of anthropomorphism. Following the experiment, feedback from users based on the lowest scores of Group 4 suggested that they found the behavior and personality change process of real humans too stable, with barely any changes observable. This points to a potential issue in experimental design: real humans indeed do not exhibit many behavioral or personality changes over short periods and our routines might be somewhat mechanical and dull in a sense. However, overall, the comparison of the first three groups in the ablation experiment, combined with the user credibility rating, effectively supports the three research hypotheses focused on in the study: 1) Evolving Agents can construct agents with perceptible and believable differentiated personality traits; 2) In the interactive simulation process, the personalities of Evolving Agents can undergo perceptible and believable evolution; 3) The personality evolution of Evolving Agents can influence their behavior patterns to undergo perceptible and believable changes, and the ablation experiment method confirmed that Evolving Agents plays a key role in supporting these research hypotheses.
@@ 7. SUPPLEMENT INSIGHT WORKSHOP
During the inspiration experiment phase, we primarily evaluated the inspirational effects and limitations of Evolving Agents as design probes on designers’ thinking and design processes. We recruited 9 users from the design field, aged between 22-50 years, covering multiple design directions such as architectural design, industrial design, interaction design, game design, etc., all with at least a bachelor’s degree.
In the experiment, we first had the designer participants spend 30 minutes observing the daily logs of personality and behavior for 3 agents, discussing these logs, and writing down design insights generated from them on paper. Following that, we facilitated a brainstorming session where each designer discussed and associated their own and others’ insights, continuously adding new insights.
Lastly, we asked each designer to step forward to present their summary of views and fill out a five-point scale questionnaire on the inspirational effect of Evolving Agents as design probes. The scoring results are as shown in the following graph.
Most designers were surprised by such a system design and provided us with many innovative ideas on how such embodied agent simulation systems could be integrated with design processes. For example, a discussant with a background in environmental design believed that this system could be used to simulate users’ experiences and feedback in certain spatial environments, thus obtaining firsthand design opinions. A student specializing in drama and film thought it could be possible to draw inspiration from the Agents’ lives in a sandbox world to aid in their story creation, while a student of service design believed this system could help explore new service scenarios tailored to specific groups.
@@ 8.1. Interdisciplinary Perspectives on Agent architecture Construction Thinking
Currently, there are many simulation Agent architectures with relatively complete capabilities in behavior, learning and interaction. The architecture of these architectures essentially continues the traditional design philosophy of agents, built around a machine architecture concept that encompasses input, processing, and output from Perception to Mind-like thinking and then to Action. With the development of large language models to this day, their emergent general abilities and intent recognition greatly facilitate interdisciplinary theoretical transfer for researchers designing AI systems or Human-Agent Interaction (HAI), focusing on the question of ”how to create embodied agents with diverse and dynamic personalities.”
In the era of natural language models, AI can exhibit stronger transfer abilities and general capabilities, and can interact with humans through natural language. This characteristic endows natural language models with the ability to ”think about problems through natural language.” This feature allows us to imitate the entire human cognitive process, simulating human thinking and output in various modules in a natural language manner.
Based on such technological transformation, we believe that the current approach to constructing agent architectures can be divided into two steps. Firstly, map the analysis architecture used in psychology for such human behaviors into computationally feasible processes and architectures. Secondly, in the specific module design, based on relevant psychological theories and combined with human analysis and processing experience, write and debug prompt structures that the natural language models can understand. This ensures that the entire process and specific processing steps are more human-like for large language models.
@@ 8.2.1. The Prospects of Embodied Agents in Sandbox Simulation Systems
During the actual experiments, many human evaluators are confused, feeling that the experiences from all four experimental groups are quite human-like, making it difficult to clearly distinguish which one is a real human. Surprisingly, the performance of real humans is actually the lowest (possibly because our human actor has a stable mood and a rather mundane life). Most experimenters tell us they mainly rely on language style and subtle emotional differences to judge whether an agent is human or AI, but the overall feeling is already very close, reaching a degree of verisimilitude.
Furthermore, users interact with embodied agent simulation systems primarily through natural language, a mode of interaction that is very friendly to researchers from all fields, especially non-expert users, as they can easily observe and study an agent’s personality, thoughts, and behaviors directly. We believe the outputs from each module in this process can greatly aid research in design studies, psychology, sociology, anthropology, and other disciplines.
Perhaps in the future, developed AI will be able to construct highly abstract, high-dimensional models of human personality development. AI can predict a person’s personality development changes more accurately and directly without the need for us to build a simulation system to continuously supply the environment and events needed for this prediction process. The ”one-click prediction” technology model may become the main technical route for our future research on individual development and even most of human behavioral psychology.
However, we believe sandbox-style simulation systems still hold very positive value. Although black-box AI models are used in the system to execute specific units such as characters’ thinking and behavior, when we look at the entire system, the journey from planning to behavior, from behavior to thinking, and finally to personality development, is actually an interpretable and traceable ”white-box mode.” Compared to letting future developed AI directly provide answers, we believe the process information and forward-deriving logic provided by the simulation system will better enhance the interpretability and transparency of HAI systems. Users can also better understand the system and the human behavioral psychology explored behind it through interaction with such a system.
@@ 8.2.2. Limitations of large model foundational capabilities on agents
During the actual design and testing process of Evolving Agents, we also discovered several limitations in the application of embodied agents. First, due to the positive bias in the content of large language models, the growth trajectory of agents simulated over time may be overly positive, which could lead to a convergence in the development direction of agents as time progresses. Secondly, the state of Agents simulated by large language models are akin to role-playing which may prevent Agents from fully understanding the complex and subtle emotions humans experience in response to real-life events and often resulting in emotions that are too positive or somewhat confusing. Finally, since real humans are more contradictory and complex, and more perplexed and repetitive, whether it is possible to simulate humans in a more complex and realistic manner may require further in-depth research by psychologists and computer scientists together.
@@ 8.3. Ethics and Societal Impact
As the technology of large language models and agent architectures evolves, this field will also trigger more reflection on ethics and morality. Firstly, during the usage process, designers and researchers from other fields need to be cautious of becoming overly reliant on embodied agents. These agents can only simulate the thinking and behavior processes of humans but cannot fully display the complexity and diversity of real humans. Due to the limitations of the architecture, the characteristics of underlying models, and the differences between human cognition and AI cognition, AI will inevitably have instances of inaccurately simulating real humans. Therefore, we suggest using embodied agents only as divergent tools in the early stages of design or research. In actual design and in-depth study, real humans should still be the subject of research to truly understand people.
Secondly, for the whole society, the increasing capabilities and human-likeness of embodied agents will make it increasingly difficult for humans to distinguish between AI and people. Humans may form broader and more profound emotional connections with AI agents, leading to changes in social relationships. This may inspire us to reflect on connections between humans and society. Although we will inevitably enter a world where virtual and reality merge, we still recommend employing more means and methods to help humans focus more on real life and establish more connections with people around them.
@@ 9. CONCLUSION
This paper introduces Evolving Agents, interactive computational agents that simulate the diverse and dynamic human personality. We describe an novel architecture based on psychological model mapping, which can perform embodied behaviors according to differentiated personalities in an interactive simulation system and undergo continuous personality evolution based on external information during the interaction process. Evaluations in the sandbox simulation system show that our architecture creates agents with believable and perceptible personalized behaviors as well as a bidirectional evolution process of personality behaviors. Additionally, through workshops, we have preliminarily verified that such systems have a significant design inspiration during the interaction process with designers, demonstrating the development potential of embodied agents. Looking ahead, we suggest that Evolving Agents can be applied in the early stages of design, as well as in related research in psychology, sociology, and anthropology.
@@ 10. ACKNOWLEDGEMENTS
This article was funded by the Important International Academic Conference Award Fund for Graduate Students of Tongji University and the Key Technology Research and Demonstration Project of Hydrogen Town Fuel Cell Integrated Energy System.

### v3

@@ 6.1. Evaluation Procedure
We recruited 30 human evaluators (14 male, 16 female) who signed informed consent forms to participate in our study. Their ages ranged from 18 to 34 years, encompassing various disciplines and professions, with good reading and analytical skills, and all having a bachelor’s degree or higher education.
Before the formal experiment, we conducted a preliminary study. In this pre-experiment, we first selected a 10-day log of one of the agents. We presented it through front-end animations on the simulation platform for two human evaluators to assess. During the pre-experiment, we identified the following issues: 1) In practice, human evaluators were overly distracted by images and other irrelevant information, making it difficult for them to focus on the Agent’s personality growth and behavioral changes. 2) The amount of textual information from 10 days of evolution was too overwhelming for general evaluators, making it difficult for them to fully understand the complete process of the Agent’s changes over the ten days; 3) The evolution of the Agent’s personality and behavior is a highly abstract and generalized simulation of human personality evolution. Observing an agent for a long simulation time is time-consuming for human evaluators. Human evaluators must detach from the normal human perception of time and focus on evaluating the evolution process.
Therefore, we used text records of an agent’s logs over three consecutive days for user assessment in the formal experiment. Additionally, before the experiment, we emphasized to the human evaluators the importance of reducing consideration of the time factor and focusing on the believability of the evolutionary process itself.
In the formal experiment, we first asked human evaluators to spend about an hour reading and taking notes on the logs of three agents and four ablated versions of one Agent. We also asked them to use the think-aloud method to communicate their thoughts to the experimenters during the reading process. After completing the reading, we asked them to complete questionnaires for the control and ablation experiments based on their notes. During this process, they could continually refer to the reading materials to ensure the accuracy of their choices. The questionnaires we set up focused on two main aspects: user believability rating and ablation control experiments.
@@ 6.3. Ablation Experiment
To validate the positive impact of various modules within the Evolving Agents on personality differentiation and personality evolution in terms of perceptibility and believability, we conducted two degrees of ablation studies on the architecture of the Evolving Agent. Compared to the full Evolving Agents architecture, Group 2 had the Cognitive Feelings, Insight, and Character Growth parts of the Personality system removed, consistent with section 5.1. At the same time, Group 3 was further simplified from Group 2 by simplifying the Character Structure, adopting a character description method similar to those used in other existing embodied agent sandbox simulation systems. We asked human evaluators to rank the human-likeness of the growth trajectories of personality and behavior for four experimental groups and to describe the reasons for their rankings in words.
@@
We obtained 30 sets of ranking data from user studies in the ablation experiment. Based on this data, we calculated the TrueSkill ranking scores for the four experimental groups and computed Cohen’s d values to quantify the strength of differences between groups. In the TrueSkill scoring system, the $\mu$ value represents the average performance level of an experimental group, while the $\sigma$ value measures the uncertainty of the scores. Additionally, to explore the statistical significance of these results, we conducted a Kruskal-Wallis test, a non-parametric alternative to one-way ANOVA, to confirm the overall statistical significance of differences under different experimental groups. To pinpoint the exact sources of differences, namely which specific group comparisons showed significant differences, we used the Dunn test for pairwise comparisons between experimental groups, and p-values were adjusted using the Holm-Bonferroni method. The experimental results are illustrated in a bar chart, with the height of the bars corresponding to the $\mu$  value and the top of the bars marked with error lines, indicating the $\sigma$  value.
@@
The final Result shows, Group 1 achieved the best scoring results ($\mu$=29.89; $\sigma$=1.61), followed by Group 2 ($\mu$=23.18; $\sigma$=1.13), and then Group 3 ($\mu$=22.95; $\sigma$=1.15). Comparing the highest-scoring Group 1 with the composite scores of the other three groups, Cohen’s d value of 5.33 indicates that Group 1 performed significantly better than Groups 2, 3, and 4 combined. The Dunn posthoc test, both uncorrected and corrected using the Holm-Bonferroni method, showed that the difference between Group 1 and all other groups (Groups 2, 3, and 4) remained significant. At the same time, there was no significant difference among Groups 2, 3, and 4.
@@
Among all experimental control groups, Group 4 performed the worst ($\mu$=21.16; $\sigma$=1.18 ). Surprisingly, users perceived the human group to have the lowest level of anthropomorphism. Following the experiment, feedback from users based on the lowest scores of Group 4 suggested that they found the behavior and personality change process of real humans too stable, with barely any changes observable. This points to a potential issue in experimental design: real humans, indeed do not exhibit many behavioral or personality changes over short periods, and our routines might be somewhat mechanical and dull in a sense. However, overall, the comparison of the first three groups in the ablation experiment, combined with the user credibility rating, effectively supports the three research hypotheses focused on in the study: 1) Evolving Agents can construct agents with perceptible and believable differentiated personality traits; 2) In the interactive simulation process, the personalities of Evolving Agents can undergo perceptible and believable evolution; 3) The personality evolution of Evolving Agents can influence their behavior patterns to undergo perceptible and believable changes, and the ablation experiment method confirmed that Evolving Agents plays a key role in supporting these research hypotheses.
@@ 7. SUPPLEMENT INSIGHT WORKSHOP
During the inspiration experiment phase, we primarily evaluated Evolving Agents’ inspirational effects and limitations as design probes on designers’ thinking and design processes. We recruited nine users from the design field, aged between 22-50 years, covering multiple design directions such as architectural design, industrial design, interaction design, game design, etc., all with at least a bachelor’s degree.
In the experiment, we first had the designer participants spend 30 minutes observing the daily logs of personality and behavior for three agents, discussing them, and writing down design insights generated from them on paper. We then facilitated a brainstorming session where each designer discussed and associated their own and others’ insights, continuously adding new insights.
Lastly, we asked each designer to step forward to present their summary of views and fill out a five-point scale questionnaire on the inspirational effect of Evolving Agents as design probes. The scoring results are shown in the following graph.
Most designers were surprised by such a system design and provided us with many innovative ideas on how such embodied agent simulation systems could be integrated with design processes. For example, a discussant with a background in environmental design believed that this system could simulate users’ experiences and feedback in certain spatial environments, thus obtaining firsthand design opinions. A student specializing in drama and film thought it could be possible to draw inspiration from the Agents’ lives in a sandbox world to aid in their story creation. In contrast, a student of service design believed this system could help explore new service scenarios tailored to specific groups.
@@ 8.1. Interdisciplinary Perspectives on Agent Architecture Construction Thinking
Currently, many simulation Agent architectures have relatively complete capabilities in behavior, learning, and interaction. The architecture of these architectures essentially continues the traditional design philosophy of agents, built around a machine architecture concept that encompasses input, processing, and output from Perception to Mind-like thinking and then to Action. With the development of large language models to this day, their emergent general abilities and intent recognition greatly facilitate interdisciplinary theoretical transfer for researchers designing AI systems or Human-Agent Interaction (HAI), focusing on ”how to create embodied agents with diverse and dynamic personalities.”
In the era of natural language models, AI can exhibit stronger transfer abilities and general capabilities and can interact with humans through natural language. This characteristic endows natural language models with the ability to ”think about problems through natural language.” This feature allows us to imitate the entire human cognitive process, simulating human thinking and output in various modules in a natural language manner.
Based on such technological transformation, we believe the current approach to constructing agent architectures can be divided into two steps. First, map the psychology analysis architecture for such human behaviors into computationally feasible processes and architectures. Secondly, in the specific module design, based on relevant psychological theories and combined with human analysis and processing experience, write and debug prompt structures that the natural language models can understand. This ensures that the entire process and specific processing steps are more human-like for large language models.
@@ 8.2.1. The Prospects of Embodied Agents in Sandbox Simulation Systems
During the actual experiments, many human evaluators are confused, feeling that the experiences from all four experimental groups are quite human-like, making it difficult to distinguish which one is a real human clearly. Surprisingly, the performance of real humans is the lowest (possibly because our human actor has a stable mood and a rather mundane life). Most experimenters tell us they mainly rely on language style and subtle emotional differences to judge whether an agent is human or AI. Still, the overall feeling is already very close to reaching a degree of authenticity.
Furthermore, users interact with embodied agent simulation systems primarily through natural language, which is very friendly to researchers from all fields, especially non-expert users, as they can easily observe and study an agent’s personality, thoughts, and behaviors directly. We believe the outputs from each module in this process can greatly aid research in design studies, psychology, sociology, anthropology, and other disciplines.
Perhaps in the future, developed AI can construct highly abstract, high-dimensional models of human personality development. AI can predict a person’s personality development changes more accurately and directly without the need for us to build a simulation system to continuously supply the environment and events needed for this prediction process. The ”one-click prediction” technology model may become the main technical route for future research on individual development and even most human behavioral psychology.
However, we believe sandbox-style simulation systems still hold very positive value. Although black-box AI models are used in the system to execute specific units such as characters’ thinking and behavior, when we look at the entire system, the journey from planning to behavior, from behavior to thinking, and finally to personality development, is an interpretable and traceable ”white-box mode.” Compared to letting future-developed AI directly provide answers, we believe the process information and forward-deriving logic provided by the simulation system will better enhance the interpretability and transparency of HAI systems. Users can also better understand the system and the human behavioral psychology explored behind it through interaction with such a system.
@@ 8.2.2. Limitations of large model foundational capabilities on agents
During the actual design and testing of Evolving Agents, we also discovered several limitations in applying embodied agents. First, due to the positive bias in the content of large language models, the growth trajectory of agents simulated over time may be overly positive, which could lead to a convergence in the development direction of agents as time progresses. Secondly, the state of Agents simulated by large language models is akin to role-playing, which may prevent Agents from fully understanding the complex and subtle emotions humans experience in response to real-life events, often resulting in emotions that are too positive or somewhat confusing. Finally, since real humans are more contradictory, complex, perplexed, and repetitive, whether it is possible to simulate humans in a more complex and realistic manner may require further in-depth research by psychologists and computer scientists.
@@ 8.3. Ethics and Societal Impact
As the technology of large language models and agent architectures evolves, this field will also trigger more reflection on ethics and morality. Firstly, designers and researchers from other fields must be cautious of becoming overly reliant on embodied agents during the usage process. These agents can only simulate humans’ thinking and behavior processes but cannot fully display the complexity and diversity of real humans. Due to the limitations of the architecture, the characteristics of underlying models, and the differences between human cognition and AI cognition, AI will inevitably have instances of inaccurately simulating real humans. Therefore, we suggest using embodied agents only as divergent tools in the early stages of design or research. In actual design and in-depth study, real humans should still be the subject of research to truly understand people.
Secondly, for the whole society, the increasing capabilities and human-likeness of embodied agents will make it increasingly difficult for humans to distinguish between AI and people. Humans may form broader and more profound emotional connections with AI agents, leading to changes in social relationships. This may inspire us to reflect on the connections between humans and society. Although we will inevitably enter a world where virtual and reality merge, we still recommend employing more means and methods to help humans focus more on real life and establish more connections with people around them.
@@ 9. CONCLUSION
This paper introduces Evolving Agents, interactive computational agents that simulate the diverse and dynamic human personality. We describe a novel architecture based on psychological model mapping, which can perform embodied behaviors according to differentiated personalities in an interactive simulation system and undergo continuous personality evolution based on external information during the interaction process. Evaluations in the sandbox simulation system show that our architecture creates agents with believable and perceptible personalized behaviors and a bidirectional evolution process of personality behaviors. Additionally, through workshops, we have preliminarily verified that such systems have a significant design inspiration during interaction with designers, demonstrating the development potential of embodied agents. Looking ahead, we suggest that Evolving Agents can be applied in the early stages of design and related research in psychology, sociology, and anthropology.
@@ 10. ACKNOWLEDGEMENTS
This article was funded by the Important International Academic Conference Award Fund for Graduate Students of Tongji University and the Key Technology Research and Demonstration Project of Hydrogen Town Fuel Cell Integrated Energy System(allportPatternGrowthPersonality1961).